# Analog Computing for Signal Processing and Communications – Part I: Computing with Microwave Networks

### v2

@@ III Analysis of a Microwave Linear Analog Computer
We have introduced and modeled the concept of MiLAC, highlighting that it consists of a reconfigurable microwave network with multiple ports where input signals are applied and output signals are read. In this section, we characterize the output $\mathbf{v}$ as a function of the input $\mathbf{u}$ and the admittance matrix of the microwave network $\mathbf{Y}$. For better clarity, we distinguish between the cases where the input is applied only to some ports ($N<P$) and to all ports ($N=P$).
@@ III-A Analysis of a MiLAC with Input on Some Ports
The proposition follows from [38, Theorem 2.1]. ∎
@@
Remarkably, (27)-(28) and (29)-(30) contain the expressions of the output vectors $\mathbf{v}_{1}$ and $\mathbf{v}_{2}$ as a function of the input $\mathbf{u}$ and the blocks of $\mathbf{P}$, which are directly related to the blocks of $\mathbf{Y}$ by (26), when $M>0$.
@@
Note that (27)-(28) and (29)-(30) would require computationally expensive matrix-matrix products and matrix inversion operations if computed through a digital computer. However, they can be quickly computed by a MiLAC as the EM signal propagates within the microwave network at light speed. In other words, after properly setting the input $\mathbf{u}$ and the tunable components $\{Y_{i,k}\}_{i,k=1}^{P}$, (27)-(28) and (29)-(30) can be computed with complexity $\mathcal{O}(1)$, independently from the input size $N$ and output size $P=N+M$. Remarkably, the expressions in (27)-(28) and (29)-(30) are particularly relevant for computing the LMMSE estimator and the covariance matrix of its error, as it will be clarified in Section IV.
Figure 4: Representation of an $N$-port MiLAC with input on all ports.
@@
$\mathbf{v}_{2}=\left(\left(\mathbf{P}_{21}\mathbf{P}_{11}^{-1}\mathbf{P}_{12}-\mathbf{P}_{22}\right)^{-1}\mathbf{P}_{21}\mathbf{P}_{11}^{-1}\right)\mathbf{u}$ $\hat{\mathbf{x}}_{\textnormal{LMMSE},1}=\left(\left(\mathbf{H}^{H}\mathbf{C}_{\mathbf{n}}^{-1}\mathbf{H}+\mathbf{C}_{\mathbf{x}}^{-1}\right)^{-1}\mathbf{H}^{H}\mathbf{C}_{\mathbf{n}}^{-1}\right)\mathbf{y}$ [±$\begin{bmatrix}\pm\mathbf{C}_{\mathbf{n}}&\mathbf{H}\\ \mathbf{H}^{H}&\mp\mathbf{C}_{\mathbf{x}}^{-1}\end{bmatrix}$
$\mathbf{v}_{2}=\left(\mathbf{P}_{22}^{-1}\mathbf{P}_{21}\left(\mathbf{P}_{12}\mathbf{P}_{22}^{-1}\mathbf{P}_{21}-\mathbf{P}_{11}\right)^{-1}\right)\mathbf{u}$ $\hat{\mathbf{x}}_{\textnormal{LMMSE},2}=\left(\mathbf{C}_{\mathbf{x}}\mathbf{H}^{H}\left(\mathbf{H}\mathbf{C}_{\mathbf{x}}\mathbf{H}^{H}+\mathbf{C}_{\mathbf{n}}\right)^{-1}\right)\mathbf{y}$ [±$\begin{bmatrix}\pm\mathbf{C}_{\mathbf{n}}&\mathbf{H}\\ \mathbf{H}^{H}&\mp\mathbf{C}_{\mathbf{x}}^{-1}\end{bmatrix}$
@@ III-C Primitive Operations of a MiLAC
gives in $\mathbf{v}_{2}$ the output on the last $M$ ports of the MiLAC, with $M=P-N$. This primitive is mathematically given by (28) or (30) depending on the invertibility of $\mathbf{P}_{11}$ and $\mathbf{P}_{22}$. Last, the third primitive is denoted as
@@
and returns the output on all the $P$ ports of the MiLAC, i.e., the concatenation of (39) and (40). The primitive (41) coincides with (39) when the input is applied on all the ports, i.e., $N=P$ and $M=0$. The sequential use of different primitive operations computed in the analog domain can enable the efficient analog computation of complex algorithms. Nevertheless, it remains less investigated in literature [4]-[36]. In the following, we show how to exploit these three primitives together with low-complexity digital operations to efficiently compute functions with practical interest, namely the LMMSE estimator, the matrix inversion, and the Kalman filter.
@@ IV-A LMMSE Estimator for Linear Observation Processes
where $\mathbf{H}\in\mathbb{C}^{Y\times X}$ is a known constant matrix, $\mathbf{x}\in\mathbb{C}^{X\times 1}$ is the unknown random vector with mean $\bar{\mathbf{x}}=\mathbf{0}_{X\times 1}$ and covariance matrix $\mathbf{C}_{\mathbf{x}}$, and $\mathbf{n}\in\mathbb{C}^{Y\times 1}$ is the random noise vector with mean $\bar{\mathbf{n}}=\mathbf{0}_{Y\times 1}$ and covariance matrix $\mathbf{C}_{\mathbf{n}}$. We assume that $\mathbf{C}_{\mathbf{x}}$ and $\mathbf{C}_{\mathbf{n}}$ are known and invertible, and that the cross-covariance matrix of $\mathbf{x}$ and $\mathbf{n}$ is $\mathbf{C}_{\mathbf{x}\mathbf{n}}=\mathbf{0}_{X\times Y}$. According to estimation theory, an estimator $\hat{\mathbf{x}}\in\mathbb{C}^{X\times 1}$ of $\mathbf{x}$ is any function of $\mathbf{y}$, its estimation error vector writes as $\mathbf{e}=\hat{\mathbf{x}}-\mathbf{x}$, and the mean square error (MSE) is given by $\text{E}[\|\hat{\mathbf{x}}-\mathbf{x}\|_{2}^{2}]$ [39, Chapter 2]. The LMMSE estimator $\hat{\mathbf{x}}_{\text{LMMSE}}$ of $\mathbf{x}$ is defined as the linear estimator minimizing the MSE, i.e.,
@@
The proposition follows from [39, Theorem 12.1]. ∎
@@ IV-B Computing the LMMSE Estimator with a MiLAC
are two values of $\mathbf{P}_{\text{LMMSE}}$ giving the desired output. Note that this MiLAC has $N=Y$ input ports and $M=X$ output ports with no input, giving a total of $P=X+Y$ ports. The steps necessary to compute the LMMSE estimator in the analog domain with a MiLAC are summarized in Alg. 1. Note that analog computers have been proposed to compute generalized regression, ridge regression, and linear regression through resistive memory arrays [29, 30, 35]. Interestingly, all these operations can be seen as special cases of the LMMSE estimator, confirming the generality of our model.
1:$\mathbf{y}$, $\mathbf{H}$, $\mathbf{C}_{\mathbf{x}}^{-1}$, and $\mathbf{C}_{\mathbf{n}}$.
2:$\hat{\mathbf{x}}_{\text{LMMSE}}$.
@@
4:$\hat{\mathbf{x}}_{\text{LMMSE}}=\text{MiLAC}_{2}(\mathbf{y},\{Y_{i,k}\}_{i,k=1}^{P})$.
@@
1:$\mathbf{H}$, $\mathbf{C}_{\mathbf{x}}^{-1}$, and $\mathbf{C}_{\mathbf{n}}$.
2:$\mathbf{C}_{\mathbf{e}}$.
@@
Remark 3.
A MiLAC can compute the LMMSE estimator with significantly reduced computational complexity compared to a digital computer, as no matrix-matrix product or matrix inversion operation is required. While a MiLAC computes the LMMSE estimator entirely in the analog domain requiring no digital operation, computing the values of its tunable admittance components requires a digital computer. Thus, the computational complexity of computing the LMMSE estimator with a MiLAC is given by the complexity of digitally computing the tunable admittance components in (32). By exploiting the fact that $\mathbf{P}$ in (48) is Hermitian and precomputing offline the computations involving $\mathbf{C}_{\mathbf{x}}$ and $\mathbf{C}_{\mathbf{n}}$, (32) require approximately $2XY$ and $4XY$ real operations to be digitally computed for the two cases $i\neq k$ and $i=k$, respectively, giving $6XY$ real operations in total. Conversely, the number of real operations required to compute (44) and (45) through digital computing is $38XY^{2}+8X^{2}Y+8X^{3}/$ 3 and $38X^{2}Y+8XY^{2}+8Y^{3}/$ 3, respectively, due to the two matrix-matrix products and the matrix inversion, whose complexity is assessed in the Appendix. As (44) and (45) are equivalent, the complexity of computing the LMMSE estimator digitally is $\mathrm{min}\{8XY^{2}+8X^{2}Y+8X^{3}/3,8X^{2}Y+8XY^{2}+8Y^{3}/3\}$, i.e., $8(XY^{2}+X^{2}Y+\mathrm{min}\{X^{3},Y^{3}\}/3)$. The complexity of computing the LMMSE estimator with a MiLAC and by digital computing are compared in Fig. 5, showing a gain of $2.5\times 10^{4}$ times in the case $X=Y=8192$.
@@ IV-D Inverting a Matrix with a MiLAC
The computational complexity of this matrix inversion approach is driven by the number of operations required to compute the tunable admittance components according to (38), given by approximately $4N^{2}$ real operations in total. Once the tunable admittance components are reconfigured, a MiLAC can compute the matrix inverse with only $N$ measurements, where each measurement returns a column of the matrix inverse. Thus, a MiLAC can compute the inverse of an arbitrary $N\times N$ matrix with complexity $\mathcal{O}(N^{2})$, which is significantly less than the complexity of matrix inversion with digital computing given by $38N^{3}/$ 3, as discussed in the Appendix. The complexity of inverting a matrix with a MiLAC and by digital computing are compared in Fig. 6, showing a gain of $5.5\times 10^{3}$ times in the case $N=8192$.
@@ V-A Kalman Filter
Consider a discrete-time dynamical process whose state at time $t$, denoted as $\mathbf{x}_{t}\in\mathbb{C}^{X\times 1}$, is a linear function of the state at time $t-1$, denoted as $\mathbf{x}_{t-1}\in\mathbb{C}^{X\times 1}$, given by
@@
where $\mathbf{H}_{t}\in\mathbb{C}^{Y\times X}$ is a known observation matrix and $\mathbf{n}_{t}\in\mathbb{C}^{Y\times 1}$ is the random observation noise with mean $\bar{\mathbf{n}}=\mathbf{0}_{Y\times 1}$ and covariance matrix $\mathbf{N}$ [^2] For simplicity, we use $X$ and $Y$ to denote the dimensionalities in both the LMMSE estimator and the Kalman filter, relying on context to distinguish their meaning.. The goal of the Kalman filter is to provide at time $t$ an estimate of the state $\mathbf{x}_{t}$ as a function of the estimate of the state $\mathbf{x}_{t-1}$ at time $t-1$ and the noisy observation $\mathbf{y}_{t}$ [39, Chapter 13], [41]. To this end, the Kalman filter operates in two phases, namely “prediction” and “correction” [39, Chapter 13], [41].
@@
as a function of the a posteriori estimate error covariance matrix at time $t-1$ denoted as $\mathbf{R}_{t-1\mid t-1}$.
1:$\mathbf{A}_{t}$, $\mathbf{H}_{t}$, $\mathbf{M}$, $\mathbf{N}$, $\hat{\mathbf{x}}_{t-1\mid t-1}$, and $\mathbf{R}_{t-1\mid t-1}$.
2:$\hat{\mathbf{x}}_{t\mid t}$ and $\mathbf{R}_{t\mid t}$.
3:Prediction:
4:$\hat{\mathbf{x}}_{t\mid t-1}=\mathbf{A}_{t}{\hat{\mathbf{x}}}_{t-1\mid t-1}$.
5:$\mathbf{R}_{t\mid t-1}=\mathbf{A}_{t}\mathbf{R}_{t-1\mid t-1}\mathbf{A}_{t}^{H}+\mathbf{M}$.
6:Correction:
7:$\mathbf{K}_{t}=\mathbf{R}_{t\mid t-1}\mathbf{H}_{t}^{H}\left(\mathbf{H}_{t}{\mathbf{R}}_{t\mid t-1}\mathbf{H}_{t}^{H}+\mathbf{N}\right)^{-1}$.
8:$\hat{\mathbf{x}}_{t\mid t}=\hat{\mathbf{x}}_{t\mid t-1}+\mathbf{K}_{t}\left(\mathbf{y}_{t}-\mathbf{H}_{t}\hat{\mathbf{x}}_{t\mid t-1}\right)$.
9:$\mathbf{R}_{t\mid t}=\left(\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t}\right)\mathbf{R}_{t\mid t-1}$.
Algorithm 4 Kalman filter
1:$\mathbf{A}_{t}$, $\mathbf{H}_{t}$, $\mathbf{M}$, $\mathbf{N}$, $\hat{\mathbf{x}}_{t-1\mid t-1}$, and $\mathbf{R}_{t-1\mid t-1}^{-1}$.
2:$\hat{\mathbf{x}}_{t\mid t}$ and $\mathbf{R}_{t\mid t}^{-1}$.
3:Prediction:
4:$\hat{\mathbf{x}}_{t\mid t-1}=\mathbf{A}_{t}{\hat{\mathbf{x}}}_{t-1\mid t-1}$.
6:Correction:
7:$\bar{\mathbf{y}}_{t}=\mathbf{y}_{t}-\mathbf{H}_{t}\hat{\mathbf{x}}_{t\mid t-1}$.
9:$\hat{\mathbf{x}}_{t\mid t}=\bar{\mathbf{x}}_{t\mid t}+\hat{\mathbf{x}}_{t\mid t-1}$.
11:Set $\mathbf{R}_{t\mid t}^{-1}$ by Alg. 3 with Input: $\mathbf{R}_{t\mid t}$.
Algorithm 5 Kalman filter with MiLAC
@@
refining the a priori estimate error covariance matrix $\mathbf{R}_{t\mid t-1}$. The two phases “prediction” and “correction” are sequentially repeated at each time step, as reported in Alg. 4333The provided complex-valued Kalman filter implicitly assumes circularly symmetric initial state, state noise, and observation noise. To treat non-circularly symmetric random variables, augmented complex-valued Kalman filters have been proposed [42].. This process ensures that the Kalman filter output $\hat{\mathbf{x}}_{t\mid t}$ is the LMMSE estimator of $\mathbf{x}_{t}$ [39, Chapter 13].

### v3

@@ III Analysis of a Microwave Linear Analog Computer
We have introduced and modeled the concept of MiLAC, highlighting that it consists of a reconfigurable microwave network with multiple ports where input signals are applied and output signals are read. In this section, we characterize the output $\mathbf{v}$ as a function of the input $\mathbf{u}$ and the admittance matrix of the microwave network $\mathbf{Y}$.  [^3] Fully-connected microwave networks have also been studied to implement beyond-diagonal RIS (BD-RIS) [45]. However, the analysis of this section differs from previous literature since a MiLAC receives (and returns) input (and output) signals directly at its ports, while the signals at the ports of a BD-RIS are typically not considered in a wireless system. For better clarity, we distinguish between the cases where the input is applied only to some ports ($N<P$) and to all ports ($N=P$).
@@ III-A Analysis of a MiLAC with Input on Some Ports
The proposition follows from [46, Theorem 2.1]. ∎
@@
Remarkably, (27)-(28) and (29)-(30) contain the expressions of the output vectors $\mathbf{v}_{1}$ and $\mathbf{v}_{2}$ as a function of the input $\mathbf{u}$ and the blocks of $\mathbf{P}$, which are directly related to the blocks of $\mathbf{Y}$ by (26), when $M>0$. The expressions of $\mathbf{v}_{1}$ and $\mathbf{v}_{2}$ show that a MiLAC can perform non-linear computations, such as matrix inversions, while being implemented with a linear microwave network. This is possible because the input data, i.e., $\mathbf{P}$ and $\mathbf{u}$, is partially encoded into the MiLAC structure (specifically, the blocks of $\mathbf{P}$ are encoded into the tunable admittance components), and not only into the input voltage $\mathbf{u}$. The non-linear mapping between the structure of a linear microwave network and its transfer function has recently received significant attention under the term “structural non-linearity” to implement non-linear physical NNs based on linear wave systems [26, 47, 48].
Figure 4: Representation of an NN-port MiLAC with input on all ports.
@@
The expressions in (27)-(28) and (29)-(30) would require computationally expensive matrix-matrix products and matrix inversion operations if computed through a digital computer. However, they can be quickly computed by a MiLAC as the EM signal propagates within the microwave network at a high fraction of the light speed [49]. In other words, after properly setting the input $\mathbf{u}$ and the tunable components $\{Y_{i,k}\}_{i,k=1}^{P}$, (27)-(28) and (29)-(30) can be computed with complexity $\mathcal{O}(1)$, independently from the input size $N$ and output size $P=N+M$. Remarkably, the expressions in (27)-(28) and (29)-(30) are particularly relevant for computing the LMMSE estimator and the covariance matrix of its error, as it will be clarified in Section IV.
@@
$\mathbf{v}_{2}=\left(\left(\mathbf{P}_{21}\mathbf{P}_{11}^{-1}\mathbf{P}_{12}-\mathbf{P}_{22}\right)^{-1}\mathbf{P}_{21}\mathbf{P}_{11}^{-1}\right)\mathbf{u}$ $\hat{\mathbf{x}}_{\textnormal{LMMSE},1}=\left(\left(\mathbf{H}^{H}\mathbf{C}_{\mathbf{n}}^{-1}\mathbf{H}+\mathbf{C}_{\mathbf{x}}^{-1}\right)^{-1}\mathbf{H}^{H}\mathbf{C}_{\mathbf{n}}^{-1}\right)\mathbf{y}$ $\begin{bmatrix}\pm\mathbf{C}_{\mathbf{n}}&\mathbf{H}\\ \mathbf{H}^{H}&\mp\mathbf{C}_{\mathbf{x}}^{-1}\end{bmatrix}$
$\mathbf{v}_{2}=\left(\mathbf{P}_{22}^{-1}\mathbf{P}_{21}\left(\mathbf{P}_{12}\mathbf{P}_{22}^{-1}\mathbf{P}_{21}-\mathbf{P}_{11}\right)^{-1}\right)\mathbf{u}$ $\hat{\mathbf{x}}_{\textnormal{LMMSE},2}=\left(\mathbf{C}_{\mathbf{x}}\mathbf{H}^{H}\left(\mathbf{H}\mathbf{C}_{\mathbf{x}}\mathbf{H}^{H}+\mathbf{C}_{\mathbf{n}}\right)^{-1}\right)\mathbf{y}$ $\begin{bmatrix}\pm\mathbf{C}_{\mathbf{n}}&\mathbf{H}\\ \mathbf{H}^{H}&\mp\mathbf{C}_{\mathbf{x}}^{-1}\end{bmatrix}$
@@ III-C Primitive Operations of a MiLAC
gives in $\mathbf{v}_{2}$ the output on the last $M$ ports of the MiLAC, with M=PNM=P-N. This primitive is mathematically given by (28) or (30) depending on the invertibility of $\mathbf{P}_{11}$ and $\mathbf{P}_{22}$. Last, the third primitive is denoted as
@@
and returns the output on all the $P$ ports of the MiLAC, i.e., the concatenation of (39) and (40). The primitive (41) coincides with (39) when the input is applied on all the ports, i.e., $N=P$ and $M=0$. The sequential use of different primitive operations computed in the analog domain can enable the efficient analog computation of complex algorithms. Nevertheless, it remains less investigated in the literature [4]-[43]. In the following, we show how to exploit these three primitives together with low-complexity digital operations to efficiently compute functions with practical interest, namely the LMMSE estimator, the matrix inversion, and the Kalman filter.
@@ IV-A LMMSE Estimator for Linear Observation Processes
where $\mathbf{H}\in\mathbb{C}^{Y\times X}$ is a known constant matrix, $\mathbf{x}\in\mathbb{C}^{X\times 1}$ is the unknown random vector with mean $\bar{\mathbf{x}}=\mathbf{0}_{X\times 1}$ and covariance matrix $\mathbf{C}_{\mathbf{x}}$, and $\mathbf{n}\in\mathbb{C}^{Y\times 1}$ is the random noise vector with mean $\bar{\mathbf{n}}=\mathbf{0}_{Y\times 1}$ and covariance matrix $\mathbf{C}_{\mathbf{n}}$. We assume that $\mathbf{C}_{\mathbf{x}}$ and $\mathbf{C}_{\mathbf{n}}$ are known and invertible, and that the cross-covariance matrix of $\mathbf{x}$ and $\mathbf{n}$ is $\mathbf{C}_{\mathbf{x}\mathbf{n}}=\mathbf{0}_{X\times Y}$. According to estimation theory, an estimator $\hat{\mathbf{x}}\in\mathbb{C}^{X\times 1}$ of $\mathbf{x}$ is any function of $\mathbf{y}$, its estimation error vector writes as $\mathbf{e}=\hat{\mathbf{x}}-\mathbf{x}$, and the mean square error (MSE) is given by $\text{E}[\|\hat{\mathbf{x}}-\mathbf{x}\|_{2}^{2}]$ [50, Chapter 2]. The LMMSE estimator $\hat{\mathbf{x}}_{\text{LMMSE}}$ of $\mathbf{x}$ is defined as the linear estimator minimizing the MSE, i.e.,
@@
The proposition follows from [50, Theorem 12.1]. ∎
@@ IV-B Computing the LMMSE Estimator with a MiLAC
[…]
@@
[…]
@@
[…]
@@
are two values of $\mathbf{P}_{\text{LMMSE}}$ giving the desired output. Both values of $\mathbf{P}_{\text{LMMSE}}$ included in (48) result in the desired LMMSE estimator expression, as it can be shown by analyzing the expression of the MiLAC output in (28) and (30). The steps necessary to compute the LMMSE estimator in the analog domain with a MiLAC are summarized in Alg. 1, where the MiLAC has $N=Y$ input ports and $M=X$ output ports with no input, namely a total of $P=X+Y$ ports. Note that analog computers have been proposed to compute generalized regression, ridge regression, and linear regression through resistive memory arrays [33, 34, 42]. Interestingly, all these operations can be seen as special cases of the LMMSE estimator, confirming the generality of our model.
@@
Remark 3.
A MiLAC can compute the LMMSE estimator with significantly reduced computational complexity compared to a digital computer, as no matrix-matrix product or matrix inversion operation is required. While a MiLAC computes the LMMSE estimator entirely in the analog domain requiring no digital operation, computing the values of its tunable admittance components requires a digital computer. Thus, the computational complexity of computing the LMMSE estimator with a MiLAC is given by the complexity of digitally computing the tunable admittance components in (32), in terms of the number of arithmetic real operations required. By exploiting the fact that $\mathbf{P}$ in (48) is Hermitian and precomputing offline the computations involving $\mathbf{C}_{\mathbf{x}}$ and $\mathbf{C}_{\mathbf{n}}$, (32) require approximately $2XY$ and $4XY$ real operations to be digitally computed for the two cases $i\neq k$ and $i=k$, respectively, giving $6XY$ real operations in total. Conversely, the number of real operations required to compute (44) and (45) through digital computing is $8XY^{2}+8X^{2}Y+8X^{3}/3$ and $8X^{2}Y+8XY^{2}+8Y^{3}/3$, respectively, due to the two matrix-matrix products and the matrix inversion, whose complexity is assessed in the Appendix. As (44) and (45) are equivalent, the complexity of computing the LMMSE estimator digitally is $\mathrm{min}\{8XY^{2}+8X^{2}Y+8X^{3}/3,8X^{2}Y+8XY^{2}+8Y^{3}/3\}$, i.e., $8(XY^{2}+X^{2}Y+\mathrm{min}\{X^{3},Y^{3}\}/3)$. The complexity of computing the LMMSE estimator with a MiLAC and by digital computing are compared in Fig. 5, showing a gain of $2.5\times 10^{4}$ times in the case $X=Y=8192$.
@@ IV-D Inverting a Matrix with a MiLAC
The computational complexity of this matrix inversion approach is driven by the number of operations required to compute the tunable admittance components according to (38), given by approximately $4N^{2}$ real operations in total. Once the tunable admittance components are reconfigured, a MiLAC can compute the matrix inverse with only $N$ measurements, where each measurement returns a column of the matrix inverse. Thus, a MiLAC can compute the inverse of an arbitrary $N\times N$ matrix with complexity $\mathcal{O}(N^{2})$, which is significantly less than the complexity of matrix inversion with digital computing given by $8N^{3}/3$, as discussed in the Appendix. The complexity of inverting a matrix with a MiLAC and by digital computing are compared in Fig. 6, showing a gain of $5.5\times 10^{3}$ times in the case $N=8192$.
@@ V Analog Computing for the Kalman Filter
[…]
3:Prediction:
[…]
6:Correction:
[…]
Algorithm 4 Kalman filter
[…]
3:Prediction:
[…]
6:Correction:
[…]
11:Set $\mathbf{R}_{t\mid t}^{-1}$ by Alg. 3 with Input: $\mathbf{R}_{t\mid t}$.
Algorithm 5 Kalman filter with MiLAC
@@ V-A Kalman Filter
Consider a discrete-time dynamical process whose state at time $t$, denoted as $\mathbf{x}_{t}\in\mathbb{C}^{X\times 1}$, is a linear function of the state at time t1t-1, denoted as $\mathbf{x}_{t-1}\in\mathbb{C}^{X\times 1}$, given by
@@
where $\mathbf{H}_{t}\in\mathbb{C}^{Y\times X}$ is a known observation matrix and $\mathbf{n}_{t}\in\mathbb{C}^{Y\times 1}$ is the random observation noise with mean $\bar{\mathbf{n}}=\mathbf{0}_{Y\times 1}$ and covariance matrix $\mathbf{N}$ [^4] For simplicity, we use $X$ and $Y$ to denote the dimensionalities in both the LMMSE estimator and the Kalman filter, relying on context to distinguish their meaning.. The goal of the Kalman filter is to provide at time $t$ an estimate of the state $\mathbf{x}_{t}$ as a function of the estimate of the state $\mathbf{x}_{t-1}$ at time t1t-1 and the noisy observation $\mathbf{y}_{t}$ [50, Chapter 13], [52]. To this end, the Kalman filter operates in two phases, namely “prediction” and “correction” [50, Chapter 13], [52].
@@
as a function of the a posteriori estimate error covariance matrix at time t1t-1 denoted as $\mathbf{R}_{t-1\mid t-1}$.
@@
refining the a priori estimate error covariance matrix $\mathbf{R}_{t\mid t-1}$. The two phases “prediction” and “correction” are sequentially repeated at each time step, as reported in Alg. 4555The provided complex-valued Kalman filter implicitly assumes circularly symmetric initial state, state noise, and observation noise. To treat non-circularly symmetric random variables, augmented complex-valued Kalman filters have been proposed [53].. This process ensures that the Kalman filter output $\hat{\mathbf{x}}_{t\mid t}$ is the LMMSE estimator of $\mathbf{x}_{t}$ [50, Chapter 13].